if their is 40 square ft worth of room , how many 6 1/2in. x 6 1/2in. tiles will you need to fill 40 square ft?
Answer by mananth(16949)

Area of room = 40 sq ft
length of side of tile = 6.5 in. = 13/2 *1/12= 13/24
Area of tile = 13 /24 *13/24= 169/576
Area of floor / area of tile = number of tiles
40 / (169/576)
40 * 576/169
= 136.3 tiles = 137 tiles